OpenStudy (anonymous):

find the inverse of f(x)=-9/x+8

OpenStudy (anonymous):

Just solve for \(x\).

OpenStudy (anonymous):

Do you know how to solve for a variable?

OpenStudy (anonymous):

Also is it \[ f(x) =-\frac{ 9}{x}+8 \]Or \[ f(x) = -\frac{9}{x+8} \]

OpenStudy (anonymous):

if f(x)=−9/x+8 then f(x)^-1 = -9/(x-8) if f(x) = -9/(x+8) then f(x)^-1 = - (x+8)/x

OpenStudy (anonymous):

@kimcarlos: no it's not that . inverse function is replacing the x by y and y by x in the original function and then solve for y.

OpenStudy (anonymous):

@Hoa thats what i exactly did..even though it does actually go with the very essence of an inverse function,

OpenStudy (anonymous):

to me, when replacing x by y. I got y = (-9/x ) -8 is the inverse function. I just compared with yours and figure out that they are not the same. sorry if I am wrong. but if I am not wrong, would you please re-check the work